OpenStudy (anonymous):

if you know that 7 x 8 =56 , how can you use the Commutative (order) Property of Multiplication to find the product of 8 x 7 ?

OpenStudy (anonymous):

56

OpenStudy (pokemon23):

8*7=56

OpenStudy (hunus):

Commutative \[a*b=b*a\] Transitive \[a=b\] \[b=c\] \[a=c\] 7*8=56 7*8=8*7 8*7=56